OpenStudy (anonymous):

If the Sun's surface temperature dropped to only half of its current temperature, how much light energy would it emit each second compared to the amount it emits now? Assume the size does not change.

OpenStudy (anonymous):

\[E\propto T \]

OpenStudy (anonymous):

@electrokid I don't understand. Would that be the formula?

OpenStudy (anonymous):

that is from Boltzmann's law. use it to get \[{E_1\over E_2} ={T_1\over T_2}\]

OpenStudy (anonymous):

@electrokid Does E stand for energy and T for temperature?

OpenStudy (anonymous):

yep whenever there is a comparison, use ratios... the english "comparision" translates to "ratio" is math

OpenStudy (anonymous):

@electrokid but how do I find out the original temperature and light emission. I'm so lost. He didn't even give us this information.

OpenStudy (anonymous):

it does not matter... it is given that T2 = T1 / 2

OpenStudy (anonymous):

I'm still so lost.

OpenStudy (anonymous):

\[ {E_1\over E_2}=\frac{T_1}{T_1\over2}=T_1\times{2\over T_1}=2\\ E_1=2E_2\implies E_2={E_1\over2} \]

OpenStudy (anonymous):

so, half the energy

OpenStudy (anonymous):

@electrokid thank you!

OpenStudy (anonymous):

you are welcome ☮